### v1

@@ 3.3 Complexity analysis
In this section, the symbol $\sim$ always denotes asymptotic equivalence, and the notation $o(\cdot)$ denotes asymptotic domination, with respect to the parameter $q$. We make the following assumption, justified by Remarks 10 and 11.
@@ 3.3.1 How many collisions do we need to find?
where ${k^{\perp}}=\dim(\mathcal{C}^{\perp})=mn-k$. The number of $\mathcal{C}_{A}$ with one-dimensional hull is therefore equivalent to $q^{{k^{\perp}}-2}$ by Section 2.1. The total number of possible classes of separable characteristic polynomials is $\sim q^{m-3}$ (see Lemma 14). Thus, given a code $\mathcal{C}_{A}$, there are $\sim q^{{k^{\perp}}-m+1}$ codes $\mathcal{D}_{B}$ that yield the same characteristic polynomial. One of these is the matrix $(P^{-1})^{\top}AQ^{\top}$ we are looking for. This means that we have to list on average $\sim\max(1,q^{{k^{\perp}}-m+1})$ collisions in order to find the right one.
@@ 3.3.2 How many matrices $A$ do we need to sample in order to find enough characteristic polynomials?
Let $r=o(q^{m-3})$ be an integer. The average number of matrices to sample in Algorithm 2 in order to get $r$ distinct characteristic polynomials is $\sim qr$.
@@
This is a variant of the coupon collector’s problem. Denote by
@@
The random variable $s_{j}$ follows a geometric distribution: it is the first success of a Bernoulli variable. The parameter $p_{j}$ of this variable is computed as follows: it is the proportion, among all the elements of $(\mathcal{C}^{\perp}-\{0\})/\mathbb{F}_{q}^{\times}$, of those (equivalence classes of) matrices $A$ yielding a code $\mathcal{C}_{A}$ with one-dimensional hull and characteristic polynomial that is not among the $j$ polynomials already in the list. The number of full–rank matrices $A$ that yield a code $\mathcal{C}_{A}$ with one-dimensional hull is
@@ 3.3.3 Complexity of computing the dictionary
The average complexity of Algorithm 2 with input a $k$-dimensional code $\mathcal{C}\subset\mathbb{F}_{q}^{m\times n}$ and a desired list length $L=o(q^{m-3})$ is
$\mathcal{O}(qLk(nm^{\omega-1}+km^{2})).$
@@
In order to get $L$ distinct characteristic polynomials, Lemma 9 tells us that we need to sample $\approx qL$ matrices $A$. For each of these, we first need to compute a basis $(C_{1},\dots,C_{k})$ of $\mathcal{C}A^{\top}$, which is given by $k$ products of a matrix of size $m\times n$ by a matrix of size $n\times m$; this requires $\mathcal{O}(knm^{\omega-1})$ operations in $\mathbb{F}_{q}$. Then, we need to compute $\mathcal{C}_{A}\cap\mathcal{C}_{A}^{\perp}$, which is given by the kernel of the (symmetric) Gram matrix $(\operatorname{Tr}(C_{i}C_{j}))_{1\leqslant i,j\leqslant k}$. Computing the diagonal entries of a given product $C_{i}C_{j}$ requires $\mathcal{O}(m^{2})$ operations in $\mathbb{F}_{q}$. Hence, the Gram matrix is computed in $\mathcal{O}(k^{2}m^{2})$ operations in $\mathbb{F}_{q}$. Computing its kernel takes $\mathcal{O}(k^{\omega})$ operations in $\mathbb{F}_{q}$. When the hull has dimension 1, we then only need to compute the characteristic polynomial of the generator we have found, which is done in $\mathcal{O}(m^{\omega})$ operations [21, Thm. 1.1], and to normalize it. This normalization can be precomputed for a proportion $(1-2/q)$ of all cases (see Remark 15). So sampling $qL$ matrices to find enough hulls $h(\mathcal{C}_{A})$ takes $\mathcal{O}(qL(knm^{\omega-1}+k^{2}m^{2}+k^{\omega}+m^{\omega}))$ operations. The result follows from the fact that $k\leqslant m^{2}$ and $m\leqslant n$. ∎
Corollary 1
The average complexity of Algorithm 3 with input a $k$-dimensional code $\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$, a dictionary of length $L\geqslant q$ and a desired number $c$ of collisions is
$\mathcal{O}(cq^{m-2}k(nm^{\omega-1}+km^{2})/L).$
@@
In order to get $c$ collisions with keys of the dictionary, which are chosen uniformly within $\sim q^{m-3}$ elements, we need to construct a list of $\sim cq^{m-3}/L$ characteristic polynomials. The complexity of this was computed in the previous lemma. ∎
3.3.4 Total complexity of this reduction step
Recall that we showed earlier (see beginning of Section 3.3) that we need to find on average $\sim\max(1,q^{{k^{\perp}}-m+1})$ collisions in order to find a suitable pair $(A,B)$ such that $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$ are actually conjugate.
In the following, we will allow ourselves to replace $\mathcal{C}$ with $\mathcal{C}^{\perp}$ if needed. Recall that we need $k<m^{2}-1$ for the algorithm to work (see Remark 6). Hence, if we want to switch $\mathcal{C}$ and $\mathcal{C}^{\perp}$, we need ${k^{\perp}}<m^{2}-1$, i.e. $k>m(n-m)+1$.
Proposition 2
Suppose that either ${k^{\perp}}\leqslant k<m^{2}-1$, or $k\leqslant{k^{\perp}}<m^{2}-1$. The reduction step, i.e. running Algorithm 2 to get a dictionary of length $q^{(m-3)/2}$ and Algorithm 3 to get $\max(1,q^{\min(k,{k^{\perp}})-m+1})$ collisions takes an expected complexity of
$\mathcal{O}\left(\max(k,{k^{\perp}})(nm^{\omega-1}+\max(k,{k^{\perp}})m^{2})q^{(m-1)/2}\max(1,q^{\min(k,{k^{\perp}})-m+1})\right)$
operations in $\mathbb{F}_{q}$, and a space complexity of
$\mathcal{O}\left(m(n+m+1)q^{(m-3)/2}\right)$
elements of $\mathbb{F}_{q}$.
Proof
We may suppose ${k^{\perp}}\leqslant k$; if not, switch $\mathcal{C}$ and $\mathcal{C}^{\perp}$, $\mathcal{D}$ and $\mathcal{D}^{\perp}$. We need to find $\sim\max(1,q^{{k^{\perp}}-m+1})$ collisions among $q^{m-3}$ elements. For this, we first compute a dictionary of length $q^{(m-3)/2}$ using Algorithm 2, then find $\max(1,q^{{k^{\perp}}-m+1})$ collisions using Algorithm 3. The result is a consequence of Lemma 10 and Corollary 1. The space complexity is simple to compute: it is dominated by the number of elements of $\mathbb{F}_{q}$ in the dictionary. There are $q^{(m-3)/2}$ entries in the dictionary, and each of them contains an $m\times n$ matrix, an $m\times m$ matrix, and coefficients of a polynomial of degree $m$. ∎
Remark 12
When $n=m$, this complexity can be further reduced by considering the codes
$\mathcal{C}_{A}=\mathcal{C}A^{-1}\quad\text{for}\quad A\in\mathcal{C}\cap\operatorname{GL}_{m}(\mathbb{F}_{q}),$
instead of $\mathcal{C}A^{\top}$ for $A\in\mathcal{C}^{\perp}$. The remainder of the algorithm is unchanged. The difference is that the algorithm now only requires
$\mathcal{O}\left(\min(k,{k^{\perp}})(nm^{\omega-1}+\min(k,{k^{\perp}})m^{2})q^{(m-1)/2}\max(1,q^{\min(k,{k^{\perp}})-m+1})\right)$
operations in $\mathbb{F}_{q}$.
Remark 13
When $m=n=k$, this is even simpler, as $\min(k,k^{\perp})=k$. In that case, our algorithm requires
$\mathcal{O}\left(n^{4}q^{(n+1)/2}\right)$
operations in $\mathbb{F}_{q}$.
@@ 4 Finding the right matrix
I order to shorten the notations, we now denote by $\mathcal{C},\mathcal{D}$ the codes $\mathcal{C}_{A},\mathcal{D}_{B}$. We are in the following situation: we are given two codes $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times m}$ with one-dimensional hulls generated respectively by matrices $U,V$ and a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$. Our aim is to decide whether $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$ are conjugate by some matrix $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$, and if they are, find such a matrix.
@@
4.1 Reducing to diagonal matrices
The one-dimensional hulls of the codes $\mathcal{C}$ and $\mathcal{D}$ are respectively generated by conjugate diagonalizable matrices $U,V\in\operatorname{GL}_{m}(\mathbb{F}_{q})$. We can compute a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$, hence $h(\mathcal{D})=Rh(\mathcal{C})R^{-1}$. We are looking for a matrix $P\in\mathbb{F}_{q}^{m\times m}$ such that $\mathcal{D}=P\mathcal{C}P^{-1}$.
We know that the matrix $P\in\mathbb{F}_{q}^{m\times m}$ we are looking for satisfies $V=PUP^{-1}$. Therefore, there exists a matrix $T\in\mathbb{F}_{q}^{m\times m}$ which commutes with $U$ such that $P=RT$. Since the characteristic polynomial of $U$ is separable, we may write $T=f(U)$, where $f=\alpha_{0}+\alpha_{1}t+\dots+\alpha_{m-1}t^{m-1}\in\mathbb{F}_{q}[t]$ (see [16, Cor. IV.E.8]). We know $R$ and search for a polynomial $f$ such that
$\mathcal{D}=Rf(U)\mathcal{C}f(U)^{-1}R^{-1}.$ ($\circ$)
Under the assumption that $U,V$ have a squarefree characteristic polynomial, there is a diagonal matrix $\Delta$ and a matrix $S\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ both possibly defined over an extension $\mathbb{F}_{q^{\prime}}$ of $\mathbb{F}_{q}$ such that $V=S\Delta S^{-1}$. Such matrices $S,\Delta$ are easily computable. Therefore,
@@ 4.1 Reducing to diagonal matrices
Result: Field extension $\mathbb{F}_{q^{\prime}}$ of $\mathbb{F}_{q}$ and tuple $(\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta)$ where:
$\Delta\in\mathbb{F}_{q^{\prime}}^{m\times m}$ is diagonal and conjugate to $U,V$
$\exists f\in\mathbb{F}_{q^{\prime}}[t]$ s.t. $\mathcal{D}^{\prime}=f(\Delta)\mathcal{C}^{\prime}f(\Delta)^{-1}\subset\mathbb{F}_{q^{\prime}}^{m\times m}$
@@
return $(\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta)$
@@
4.2 Conjugating by the right matrix

### v2

@@ 3.3 Complexity analysis
In this section, the symbol $\sim$ always denotes asymptotic equivalence, and the notation $o(\cdot)$ denotes asymptotic domination, with respect to the parameter $q$. We make the following assumption, justified by Remarks 9 and 10.
@@ 3.3.1 How many matrices $A$ do we need to sample in order to find enough characteristic polynomials?
where ${k^{\perp}}=\dim(\mathcal{C}^{\perp})=mn-k$. The number of $\mathcal{C}_{A}$ with one-dimensional hull is therefore equivalent to $q^{{k^{\perp}}-2}$ by Section 2.1. The total number of possible classes of separable characteristic polynomials is $\sim q^{m-3}$ (see Lemma 15).
Thus, the dictionary constructed in Algorithm 2 will have size $L\leqslant q^{m-3}$. Then, according to the usual list collision arguments, the number $L^{\prime}$ of one-dimensional hulls to check in Algorithm 3 should satisfy
$LL^{\prime}\sim q^{{k^{\perp}}-2}.$
We have to treat two cases separately:
  1. $(i)$
    ${k^{\perp}}-2\leqslant 2(m-3)$, where the dictionary constructed by Algorithm 2 will not need to cover all the possible characteristic polynomials and we take
    $L\sim q^{\frac{{k^{\perp}}}{2}-1}\quad{\text{and}}\quad L^{\prime}\sim q^{\frac{{k^{\perp}}}{2}-1};$
  2. $(ii)$
    ${k^{\perp}}-2>2(m-3)$, where there are not enough different characteristic polynomials to have $L\sim L^{\prime}$, and we take
    $L\sim q^{m-3}\quad\text{and}\quad L^{\prime}\sim q^{{k^{\perp}}-m+1}.$
@@ 2. $(ii)$
Let $r$ be an integer. The average number of matrices to sample in Algorithm 2 in order to get $r$ distinct characteristic polynomials is $\sim qr$ if $r=o(q^{m-3})$ and $\sim qr\log(r)$ if $r\sim q^{m-3}$.
@@
The latter case is a classical result usually called coupon collector’s problem, while for the former we prove a variant. Denote by
@@
The random variable $s_{j}$ follows a geometric distribution: it is the first success of a Bernoulli variable. The parameter $p_{j}$ of this variable is computed as follows: it is the proportion, among all the elements of $(\mathcal{C}^{\perp}-\{0\})/\mathbb{F}_{q}^{\times}$, of those (equivalence classes of) matrices $A$ yielding a code $\mathcal{C}_{A}$ with one-dimensional hull and characteristic polynomial that is not among the $j$ polynomials already in the list. From Section 2.1, the number of full–rank matrices $A$ that yield a code $\mathcal{C}_{A}$ with one-dimensional hull is
@@ 3.3.2 Complexity of computing the dictionary
The average complexity of Algorithm 2 with input a $k$-dimensional code $\mathcal{C}\subset\mathbb{F}_{q}^{m\times n}$ and a desired list length $L$ is
$\mathcal{O}(qLk(nm^{\omega-1}+km^{2}))\quad$ $\text{if}\ \ L=o(q^{m-3})$
$\mathcal{O}(q^{m-2}km(nm^{\omega-1}+km^{2}))\quad$ $\text{if}\ L\sim q^{m-3}.$
@@
In order to get $L$ distinct characteristic polynomials, Lemma 9 tells us that we need to sample $\sim qL$ matrices $A$ if $L=o(q^{m-3})$ and $\sim qL\log L$ if $L\sim q^{m-3}$. In this last case, this requires to sample $\sim mq^{m-2}$ matrices. For each of these, we first need to compute a basis $(C_{1},\dots,C_{k})$ of $\mathcal{C}A^{\top}$, which is given by $k$ products of a matrix of size $m\times n$ by a matrix of size $n\times m$; this requires $\mathcal{O}(knm^{\omega-1})$ operations in $\mathbb{F}_{q}$. Then, we need to compute $\mathcal{C}_{A}\cap\mathcal{C}_{A}^{\perp}$, which is given by the kernel of the (symmetric) Gram matrix $(\operatorname{Tr}(C_{i}C_{j}))_{1\leqslant i,j\leqslant k}$. Computing the diagonal entries of a given product $C_{i}C_{j}$ requires $\mathcal{O}(m^{2})$ operations in $\mathbb{F}_{q}$. Hence, the Gram matrix is computed in $\mathcal{O}(k^{2}m^{2})$ operations in $\mathbb{F}_{q}$. Computing its kernel takes $\mathcal{O}(k^{\omega})$ operations in $\mathbb{F}_{q}$. When the hull has dimension 1, we then only need to compute the characteristic polynomial of the generator we have found, which is done in $\mathcal{O}(m^{\omega})$ operations [21, Thm. 1.1], and to normalize it. This normalization can be precomputed for a proportion $(1-2/q)$ of all cases (see Remark 13). So sampling one matrix takes $\mathcal{O}(knm^{\omega-1}+k^{2}m^{2}+k^{\omega}+m^{\omega})$ operations, which, since $k\leqslant m^{2}$ and $m\leqslant n$ gives $\mathcal{O}(k(nm^{\omega-1}+km^{2}))$. Multiplying this by the number of sampled matrices (mentioned in the beginning of the present proof) yields the result.∎
3.3.3 Total complexity of this reduction step
Recall that we allowed ourselves to replace $\mathcal{C}$ with $\mathcal{C}^{\perp}$ if needed, in order to ensure that ${k^{\perp}}\leqslant k$ and that we need $k<m^{2}-1$ for the algorithm to work (see Remark 6).
Theorem 3.1
Suppose that $m\leqslant n$ and ${k^{\perp}}\leqslant k<m^{2}-1$ and that Assumption 1 holds. Under Assumption 2, Algorithm 3 takes an expected complexity of
$\mathcal{O}\left(km(nm^{\omega-1}+km^{2})q^{\max(\frac{{k^{\perp}}}{2},{k^{\perp}}-m+2)}+m^{2\omega}q^{{k^{\perp}}-m+1}\right)$
operations in $\mathbb{F}_{q}$, and a space complexity of
$\mathcal{O}\left(({k^{\perp}}+m+1)q^{\min(\frac{{k^{\perp}}}{2}-1,m-3)}\right)$
elements of $\mathbb{F}_{q}$.
Remark 11
One can get rid of Assumption 2 by replacing the techniques of Subsection 4.1 by those of Subsection 4.2 at the cost of another sub-exponential term in the time complexity (Lemma 13). Namely, this would give an overall complexity of
$\mathcal{O}\left(km(nm^{\omega-1}+km^{2})q^{\max(\frac{{k^{\perp}}}{2},{k^{\perp}}-m+2)}+q^{{k^{\perp}}-m+1}k^{\omega-1}m^{2}q^{3\sqrt{m}/\sqrt{\log m}\log q}\right)$
operations in $\mathbb{F}_{q}$.
@@ 3.3.3 Total complexity of this reduction step
We consider separately Cases (i) and (ii) introduced in Page i. The time complexity of Algorithm 2, given by Lemma 10, is
$\mathcal{O}\left(k(nm^{\omega-1}+km^{2})q^{\max(\frac{{k^{\perp}}}{2},{k^{\perp}}-m+2)}\right).$
The cost of Algorithm 3 depends on the number of “false positives” encountered, i.e. situations where we find for $h(\mathcal{D}_{B})$ a characteristic polynomial which is an entry of the dictionary but running an Algorithm of Section 4 shows that $\mathcal{D}_{B}$ is not a conjugate of the code $\mathcal{C}_{A}$ corresponding to that entry. For each key of the dictionary, there are $\sim q^{{k^{\perp}}-2}/q^{m-3}=q^{{k^{\perp}}-m+1}$ one-dimensional hulls corresponding to this key, roughly all of which are false positives.
In Case (i), Algorithm 3 costs
$\mathcal{O}\left(km(nm^{\omega-1}+km^{2})q^{\frac{{k^{\perp}}}{2}}+m^{2\omega}q^{{k^{\perp}}-m+1}\right).$
In Case (ii), recall that the algorithms of Section 4, whose complexities are given in Lemma 11, have to be called on average $q^{{k^{\perp}}-m+1}$ times. Here, the complexity of calling the algorithms of Section 4 outweighs that of computing the dictionary, and the time complexity of Algorithm 3 is
$\mathcal{O}(m^{2\omega}q^{{k^{\perp}}-m+1}).$
The space complexity is simple to compute: it is dominated by the number $L$ of entries in the dictionary. There are $\mathcal{O}(q^{\min(\frac{{k^{\perp}}}{2}-1,m-3)})$ such entries, which consist of a matrix $A\in\mathcal{C}^{\perp}$ and a characteristic polynomial. To reduce space complexity, one may store in each of these entries only the coordinates of the matrix $A$ in a basis of $\mathcal{C}^{\perp}$ as well as the characteristic polynomial of the generator of $h(\mathcal{C}_{A})$, which amounts to ${k^{\perp}}+m+1$ field elements per entry. ∎
@@ 4 Finding the right matrix
In order to shorten the notations, we now denote by $\mathcal{C},\mathcal{D}$ the codes $\mathcal{C}_{A},\mathcal{D}_{B}$. We are in the following situation: we are given two codes $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times m}$ with one-dimensional hulls generated respectively by matrices $U,V$ and a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$. Our aim is to decide whether $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$ are conjugate by some matrix $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$, and if they are, find such a matrix.
In this section, we propose two approaches. The first one is based on multivariate polynomial system solving and has a polynomial time complexity conditioned by some assumption. The second one is based on diagonalization arguments, has a sub-exponential time complexity but does not rest on any assumption.
4.1 First approach: using polynomial system solving
The one-dimensional hulls of the codes $\mathcal{C}$ and $\mathcal{D}$ are respectively generated by conjugate matrices $U,V\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ having a squarefree characteristic polynomial. We can compute a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$, hence $h(\mathcal{D})=Rh(\mathcal{C})R^{-1}$. We are looking for a matrix $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $\mathcal{D}=P\mathcal{C}P^{-1}$. We know that the matrix $P\in\mathbb{F}_{q}^{m\times m}$ we are looking for satisfies $V=PUP^{-1}$. Therefore, there exists a matrix $T\in\mathbb{F}_{q}^{m\times m}$ which commutes with $U$ such that $P=RT$. Since the characteristic polynomial of $U$ is separable, we can write $T=f(U)$, where $f=\alpha_{0}+\alpha_{1}t+\dots+\alpha_{m-1}t^{m-1}\in\mathbb{F}_{q}[t]$ (see [16, Cor. IV.E.8]). We know $R$ and search for a polynomial $f$ such that
$\mathcal{D}=Rf(U)\mathcal{C}f(U)^{-1}R^{-1}.$ ($\circ$)
A usual argument (for instance by Cayley-Hamilton Theorem) shows that there exists a polynomial $g$ of degree $<m$ such that $f(U)^{-1}=g(U)$. Then, we can solve the following system whose unknowns are the coefficients of $f,g$:
$Rf(U)\mathcal{C}g(U)R^{-1}\subset\mathcal{D}.$ (2)
The above system is bilinear in the $m$ coefficients of $f$ and the $m$ coefficients of $g$. Thus, when linearizing, this yields $m^{2}$ unknowns while $\mathcal{C},\mathcal{D}\in\mathbb{F}_{q}^{m\times m}$ both have dimension $k$ (recall that for short we denote $\mathcal{C}_{A},\mathcal{D}_{B}$ by $\mathcal{C},\mathcal{D}$), hence the system has $k(m^{2}-k)$ equations which exceeds $m^{2}$ as soon as $11<k<m^{2}-$ 1. Thus, the linearized system is over-constrained, which encourages to make the following assumption.
Assumption 2
The linearized version of System (2) has a space of solutions of dimension $1$ when $\mathcal{C},\mathcal{D}$ are conjugate and no nonzero solution otherwise.
Lemma 11
Under Assumption 2, the calculation of the polynomial $f$ and hence of the matrix $P$ requires $\mathcal{O}(m^{2\omega})$ operations in $\mathbb{F}$.
4.2 Second approach: using diagonalization
@@
4.2.1 Reducing to diagonal matrices
Under the assumption that $U,V$ have a squarefree characteristic polynomial, there is a diagonal matrix $\Delta$ and a matrix $S\in\operatorname{GL}_{m}(\mathbb{F}_{q^{\prime}})$ both defined over an extension $\mathbb{F}_{q^{\prime}}$ of $\mathbb{F}_{q}$ such that $V=S\Delta S^{-1}$. Such matrices $S,\Delta$ are easily computable. Therefore,
@@ 4.2.1 Reducing to diagonal matrices
Result: Tuple $(\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta,S)$ where:
$S\in\operatorname{GL}_{m}(\mathbb{F}_{q^{\prime}})$ for some extension $\mathbb{F}_{q^{\prime}}/\mathbb{F}_{q}$,
$\Delta\in\operatorname{GL}_{m}(\mathbb{F}_{q^{\prime}})$ diagonal, $V=S\Delta S^{-1}$, $\mathcal{C}^{\prime}=S^{-1}R\mathcal{C}R^{-1}S$ and $\mathcal{D}^{\prime}=S^{-1}\mathcal{D}S$.
@@
return $(\mathcal{C}^{\prime},\mathcal{D}^{\prime},\Delta,S)$
@@
4.2.2 Conjugating by the right matrix